# Effects of Large Optical Depth on CD Hi  21-cm Non-Gaussian Signal

### v1

@@ Abstract
During the Cosmic Dawn (CD), the HI 21-cm optical depth ($\tau$ ) in the intergalactic medium can become significantly large. Consequently, the second and higher-order terms of $\tau$ appearing in the Taylor expansion of the HI 21-cm differential brightness temperature ($\delta T_{\rm b}$ ) become important. This introduces additional non-Gaussianity into the signal. We study the impact of large $\tau$ on statistical quantities of HI 21-cm signal using a suite of standard numerical simulations that vary X-ray heating efficiency and the minimum halo mass required to host radiation sources. We find that the higher order terms suppress statistical quantities such as skewness, power-spectrum and bispectrum. However, the effect is found to be particularly strong on the non-Gaussian signal. We find that the change in skewness can reach several hundred percent in low X-ray heating scenarios. For moderate and high X-ray heating models changes are around $\sim 40\%$ and $60\%$, respectively, for $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. This change is around $\sim 75\%$, $25\%$ and $20\%$ for low, moderate and high X-ray heating models, respectively, for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$. The change in bispectrum in both the halo cutoff mass scenarios ranges from $\sim 10\%$ to $\sim 300\%$ for low X-ray heating model. However, for moderate and high X-ray heating models the change remains between $\sim 10\%$ to $\sim 200\%$ for both equilateral and squeezed limit triangle configuration. Finally, we find that up to third orders of $\tau$ need to be retained to accurately model $\delta T_{\rm b}$, especially for capturing the non-Gaussian features in the Hi 21-cm signal.
@@
pubyear: 2025pagerange: Effects of Large Optical Depth on CD Hi  21-cm Non-Gaussian SignalEffects of Large Optical Depth on CD Hi  21-cm Non-Gaussian Signal
@@ 1 Introduction
In this paper, we study the impact of large 21-cm signal optical depth on the spatial fluctuations of the CD 21-cm signal differential brightness temperature ($\delta T_{\rm b}$). In particular, we focus on the non-Gaussian features of the CD 21-cm signal, along with its power spectrum. The 21-cm signal optical depth ($\tau$), which is proportional to the Hi density and inversely proportional to the Hi spin temperature, can become significantly large during CD, particularly in regions with high Hi density and low spin temperature. As a result, when $\tau$ is expanded in Taylor series, not only the linear term (which has been generally taken into account in the literature), but also contributions from the second and higher-order terms in the expansion of $\tau$ to the $\delta T_{\rm b}$ equation become important (see e.g., Datta et al., 2022). These higher-order terms not only change the spatial fluctuations in $\delta T_{\rm b}$, but also introduce additional non-Gaussianity in the signal. Therefore, it is important to investigate how this affects the spatial fluctuations in $\delta T_{\rm b}$, as well as the non-Gaussian properties of the signal from CD.
@@
The bispectrum has been extensively used to quantify the non-Gaussian nature of the Hi 21-cm signal from the CD-EoR (Bharadwaj & Pandey, 2005; Shimabukuro et al., 2016; Majumdar et al., 2018). The bispectrum, a Fourier conjugate of the three-point correlation function, is sensitive to correlations between different Fourier modes. Majumdar et al. (2020) presented the first comprehensive analysis of the EoR 21-cm bispectrum across all unique $k$-triangle configurations. Subsequent studies have explored various physical effects on the 21-cm bispectrum, including X-ray heating and $\rm{Ly{\alpha}}$ coupling (Watkinson et al., 2018; Kamran et al., 2021b, a, 2022), IGM physics (Kamran et al., 2021a, 2022), redshift-space distortions and light cone effect (Kamran et al., 2021b; Mondal et al., 2021; Gill et al., 2024), impact of ionizing source models and ionization morphology (Noble et al., 2024; Hutter et al., 2019), impact of distribution of neutral hydrogen islands at the end stage of reionization (Raste et al., 2024), impact of different kinds of dark matter models (Saxena et al., 2020) etc. It has also been demonstrated that the SKA-Low will be sensitive enough to detect the bispectrum signal from the reionization era (Mondal et al., 2021; Tiwari et al., 2022) and bispectrum can put tighter constraints on reionization model parameters (Tiwari et al., 2022) compared to the signal power spectrum. Furthermore, a fast and scalable estimator has been developed for computing the bispectrum of the redshifted 21 cm signal from radio interferometric observations (Gill & Bharadwaj, 2025). Efforts have also been made to directly measure the bispectrum of the signal from observational data (Trott et al., 2019).
@@
The outline of the paper is as follows. In Section 2 we will discuss about the fundamentals of the Hi 21-cm signal. We present the basic equations for computing the Hi 21-cm differential brightness temperature, both exactly and approximately as outlined in subsection 2.1. Subsection 2.2 describes the simulations and models used in this work along with the results on the global evolution of the Hi 21-cm signal. In Section 3 we discuss our results. Subsections 3.1 and 3.2 discusses the optical depth, and its distribution and the impact of large Hi optical depth on the non-Gaussian features of the 21-cm signal, focusing on the distribution of $\delta T_{\rm b}$. In subsections 3.3 and 3.4, we discussed the redshift evolution of the skewness and the impact of large optical depth on it, respectively. Subsections 3.5 and 3.6 examine the redshift evolution of the bispectrum and explores how large optical depths influence it. We also discussed the effect of large optical depth on the Hi 21-cm power spectrum in Section 4. Section 5 investigates the effect of including different orders in the optical depth expansion on bispectrum estimation. Finally, Section 6 summarizes the main results of the paper.
In this paper, we use the values of various cosmological parameters for our numerical calculations and simulations: $\Omega_{\rm b}=0.044$, $\Omega_{\rm m}=0.27$, $h=0.7$ and the present CMB temperature of $2.725$ K. These values are consistent with the WMAP results (Hinshaw et al., 2013) and lie within the error bars of the Planck results (Ade et al., 2014).
@@ 2.1 Basic equations of Hi  21-cm signal
$\tau({\bf n},z)$ $=$ $27\hskip 2.84544pt\frac{x_{\rm HI}(\textbf{n},z)}{T_{S}(\textbf{n},z)}\left[1+\delta_{\rm b}(\textbf{n},z)\right]\left(\frac{\Omega_{\rm b}h^{2}}{0.023}\right)$ (2)
$\times$ $\left(\frac{0.15}{\Omega_{\rm m}h^{2}\cdot 10}\right)^{1/2}(1+z)^{3/2}\,\mathrm{mK},$
where $x_{\rm HI}(\textbf{n},z)$ and $\delta_{\rm b}(\textbf{n},z)$ are neutral hydrogen fraction and baryonic density contrast respectively along the line of sight direction n and redshift $z$. $\Omega_{\rm b}$ and $\Omega_{\rm m}$ are the cosmological baryon density and total matter density parameters at present, respectively. $h$ is the present Hubble constant in units of $100$ km/s/Mpc. We do not consider the effects of the peculiar velocity in this study.
@@
×$\times\left(\frac{0.15}{\Omega_{\rm m}h^{2}}\frac{1+z}{10}\right)^{1/2}\left[1-\frac{T_{\gamma}(z)}{T_{\rm S}(\textbf{n},z)}\right]\,\mathrm{mK}.$ (4)
@@ 2.2 Simulations of Hi  21-cm signal during CD
We use grizzly code (Ghara et al., 2015a, b, 2018) for $21$-cm signal simulations, which is based on a one-dimensional radiative transfer scheme. This code requires a gas density field, a halo catalog and a source model as its inputs. The dark matter distribution in a cubic box of comoving length $714.29$ Mpc is obtained using an N-body code named cubep3m (Harnois-Deraps et al., 2013) which is smoothed into 600 uniform grids resulting in a spatial resolution of $1.19$ Mpc. Dark matter distributions are obtained at multiple redshifts in the range of $z=18$ to $z=10$ during CD. Dark matter halos are identified using an on-the-fly halo finder which uses the spherical overdensity method. The lowest dark matter halos identified in our simulations have masses of approximately $10^{9}$ ${\rm M}_{\odot}$ (the mass of a dark matter particle in our simulations is $\sim 10^{8}$ ${\rm M}_{\odot}$). The stellar mass available within a halo is directly related to the DM halo mass as $M_{*}=f_{*}(\frac{\Omega_{\rm b}}{\Omega_{\rm m}})M_{\rm h}$, where the star formation efficiency $f_{*}$ is assumed to be $0.03$ (Behroozi & Silk, 2015; Sun & Furlanetto, 2016). Our simulation has two different astrophysical parameters i.e., $f_{\rm x}$ and $M_{\rm h,min}$. $M_{\rm h,min}$ is the minimun mass of a dark matter halo capable of hosting sources sources that emit $\rm{Ly{\alpha}}$, X-ray and UV ionizing photons. We consider two different values of $M_{\rm h,min}$, i.e. $10^{9}$ ${\rm M}_{\odot}$ and $10^{10}$ ${\rm M}_{\odot}$, in this study. $f_{\rm x}$ is the X-ray heating efficiency parameter which is the ratio of X-ray to UV luminosity of a source. Table 1 summarizes the model parameters considered in this study. The details of these simulations are discussed in (Kamran et al., 2022).
@@
Using the DM density distribution obtained from the N-body simulation and spin temperature, neutral hydrogen distribution from the one-dimensional radiative transfer code grizzly, we generate the Hi 21-cm optical depth $\tau({\bf n},z)$ distribution and subsequently the Hi $21$-cm differential brightness temperature ($\delta T_{b}({\bf n},z)$) maps. We generate $\delta T_{b}$ maps at $23$ different redshifts covering the entire CD using both the approximated (equation (3)) and the exact (equation (1)) $\delta T_{\rm b}$ equations.
@@
We show the evolution of the global signal ($\overline{\delta T_{\rm b}}(z)$) with redshift in the upper panel of Fig. 1. We see, for all choices of $f_{\rm x}$ values and for a fixed $M_{\rm h,min}$, that the initial trend in the evolution of $\overline{\delta T_{\rm b}}(z)$ around redshift $z=18$ is very similar as only $\rm{Ly{\alpha}}$ coupling dictates the 21-cm signal during the initial phase. It helps to couple the spin temperature with the IGM kinetic temperature; therefore, the signal is seen in absorption. Later, X-ray heating becomes important and raises the IGM kinetic and spin temperature. This creates a trough-like feature in the redshift evolution of $\overline{\delta T_{\rm b}}(z)$. Further, we see that the absorption trough becomes deeper for smaller $f_{\rm x}$ values as the IGM heating starts late and the spin temperature, which is coupled to IGM kinetic temperature through $\rm{Ly{\alpha}}$ coupling, continues to decrease due to adiabatic expansion of universe. Moreover, the trough is more flattened for smaller $f_{\rm x}$ values. This is because the X-ray heating rate and cooling rate due to cosmological expansion remain similar for a longer period of time in these scenarios. As expected, higher $f_{\rm x}$ makes $\overline{\delta T_{\rm b}}(z)$ positive at higher redshift. The trough like feature in $\overline{\delta T_{\rm b}}(z)$ shifts towards lower redshifts for higher halo mass cut-off, mainly because massive DM halos are formed later and, therefore, processes such as $\rm{Ly{\alpha}}$ coupling, X-ray heating get delayed.
$f_{\rm X}$ $0.1$ $46.4$ $1000$
$M_{\rm h,min}\,(10^{9}\,{\rm M}_{\odot})$ $1.0$ $10.0$
Table 1: Different model parameters of X-ray heating efficiency $f_{\rm X}$ and minimum halo mass $M_{\rm h,min}$.
@@
From the lower panel we also see that if we vary the X-ray heating efficiency $f_{\rm x}$, both the width and depth of the troughs vary. These are maximum for $f_{\rm x}=0.1$ and minimum for $f_{\rm x}=1000$. Further, we see that $\overline{\delta T_{\rm b}}(z)$ calculated using the exact equation is smaller than the approximated method, and the difference is more for smaller values of $f_{\rm x}$. This is expected as $T_{\rm S}$ has lower values for the $f_{\rm x}=0.1$ model, i.e., $\tau$ has higher values, which leads to a greater difference in the troughs between approximated and exact $\overline{\delta T_{\rm b}}(z)$ compared to the $f_{\rm x}=1000$ model. The difference is found to be similar for different $M_{\rm h,min}$ models for a given $f_{\rm x}$ value. The percentage difference between approximated and exact $\overline{\delta T_{\rm b}}(z)$ reaches up to $6.2\%$ for [$f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$] model and $6.7\%$ for [$f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$] model. The differences reduce to roughly $5\%$ and $2.7\%$ for $f_{\rm x}=46.4$ and $1000$, respectively, for both the $M_{\rm h,min}$ values. This difference arises because of considerably large values of $\tau$ during the CD. The higher-order terms in $\exp(-\tau)$ become significant when $\tau$ is large. To elaborate on this, we plot the evolution of the mean $\tau$ with redshift in the upper panel of Fig. 1. The redshift at which the mean $\tau$ reaches its maximum corresponds to the largest difference between the exact and approximated $\overline{\delta T_{\rm b}}(z)$. The optical depth becomes maximum when the IGM kinetic and spin temperature are minimum, which occurs just before the onset of X-ray heating.

### v2

# The Effect of Large Optical Depths on the Non-Gaussian 21-cm signal from Cosmic Dawn

@@ Abstract
During the Cosmic Dawn (CD), the HI 21-cm optical depth ($\tau$ ) in the intergalactic medium can become significantly large. Consequently, the second and higher-order terms of $\tau$ appearing in the Taylor expansion of the HI 21-cm differential brightness temperature ($\delta T_{\rm b}$ ) become important. This introduces additional non-Gaussianity into the signal. We study the impact of large $\tau$ on statistical quantities of HI 21-cm signal using a suite of standard numerical simulations that vary X-ray heating efficiency and the minimum halo mass required to host radiation sources. We find that the higher order terms suppress statistical quantities such as skewness, power-spectrum and bispectrum. However, the effect is found to be particularly strong on the non-Gaussian signal. We find that the change in skewness can reach several hundred percent in low X-ray heating scenarios, whereas for moderate and high X-ray heating models the changes are around $\sim 40\%$ and $\sim 60\%$, respectively, for $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. This change is around $\sim 75\%$, $25\%$ and $20\%$ for low, moderate and high X-ray heating models, respectively, for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$. The change in bispectrum in both the halo cutoff mass scenarios ranges from $\sim 10\%$ to $\sim 300\%$ for low X-ray heating model. However, for moderate and high X-ray heating models the change remains between $\sim 10\%$ to $\sim 200\%$ for both equilateral and squeezed limit triangle configuration. Finally, we find that up to third orders of $\tau$ need to be retained to accurately model $\delta T_{\rm b}$, especially for capturing the non-Gaussian features in the Hi 21-cm signal.
@@
pubyear: 2025pagerange: The Effect of Large Optical Depths on the Non-Gaussian 21-cm signal from Cosmic DawnThe Effect of Large Optical Depths on the Non-Gaussian 21-cm signal from Cosmic Dawn
@@ 1 Introduction
In this paper, we study the impact of large 21-cm signal optical depth on the spatial fluctuations of the CD 21-cm signal differential brightness temperature ($\delta T_{\rm b}$). In particular, we focus on the non-Gaussian features of the CD 21-cm signal, along with its power spectrum. The 21-cm signal optical depth ($\tau$), which is proportional to the Hi density and inversely proportional to the Hi spin temperature, can become significantly large during CD, particularly in regions with high Hi density and low spin temperature. As a result, when $\delta T_{\rm b}$ is expanded as a Taylor series in $\tau$, not only the linear term (which has been generally taken into account in the literature), but also contributions from the second and higher-order terms in the expansion of $\delta T_{\rm b}$ become important (see e.g., Datta et al., 2022). These higher-order terms not only change the spatial fluctuations in $\delta T_{\rm b}$, but also introduce additional non-Gaussianity in the signal. Therefore, it is important to investigate how this affects the spatial fluctuations in $\delta T_{\rm b}$, as well as the non-Gaussian properties of the signal from CD.
@@
The bispectrum has been extensively used to quantify the non-Gaussian nature of the Hi 21-cm signal from the CD-EoR (Bharadwaj & Pandey, 2005; Shimabukuro et al., 2016; Majumdar et al., 2018). The bispectrum, a Fourier conjugate of the three-point correlation function, is sensitive to correlations between different Fourier modes. Majumdar et al. (2020) presented the first comprehensive analysis of the EoR 21-cm bispectrum across all unique $k$-triangle configurations. Several other studies have explored various physical effects on the 21-cm bispectrum, including X-ray heating and $\rm{Ly{\alpha}}$ coupling (Watkinson et al., 2018; Kamran et al., 2021b, a, 2022), IGM physics (Kamran et al., 2021a, 2022), redshift-space distortions and light cone effect (Kamran et al., 2021b; Mondal et al., 2021; Gill et al., 2024), impact of ionizing source models and ionization morphology (Noble et al., 2024; Hutter et al., 2019), impact of distribution of neutral hydrogen islands at the end stage of reionization (Raste et al., 2024), impact of different kinds of dark matter models (Saxena et al., 2020) etc. It has also been demonstrated that the SKA-Low will be sensitive enough to detect the 21-cm bispectrum from the reionization era (Mondal et al., 2021; Tiwari et al., 2022), and that the bispectrum can place tighter constraints on reionization model parameters compared to the HI 21-cm power spectrum (Tiwari et al., 2022; Mahida et al., 2025). Efforts have also been made to directly measure the bispectrum of the signal from observational data (Trott et al., 2019; Gill et al., 2025).
@@
The outline of the paper is as follows. In Section 2 we will discuss about the fundamentals of the Hi 21-cm signal. We present the basic equations for computing the Hi 21-cm differential brightness temperature, both exactly and approximately as outlined in subsection 2.1. Subsection 2.2 describes the simulations and models used in this work along with the results on the global evolution of the Hi 21-cm signal. In Section 3 we discuss our results. Subsections 3.1 and 3.2 discusses the optical depth, and its distribution and the impact of large Hi optical depth on the non-Gaussian features of the 21-cm signal, focusing on the distribution of $\delta T_{\rm b}$. In subsections 3.3 and 3.4, we discuss the redshift evolution of the skewness and the impact of large optical depth on it, respectively. Subsections 3.5 and 3.6 examine the redshift evolution of the bispectrum and explores how large optical depths influence it. We also discuss the effect of large optical depth on the Hi 21-cm power spectrum in Section 4. Section 5 investigates the effect of including different orders in the optical depth expansion on bispectrum estimation. Finally, Section 6 summarizes the main results of the paper.
In this paper, we use the values of various cosmological parameters for our numerical calculations and simulations: $\Omega_{\rm b}=0.044$, $\Omega_{\rm m}=0.27$, $h=0.7$ and the present Cosmic Microwave Background Radiation (CMBR) temperature of $2.725$ K. These values are consistent with the WMAP results (Hinshaw et al., 2013) and lie within the error bars of the Planck results (Ade et al., 2014).
@@ 2.1 Basic equations of Hi  21-cm signal
$\tau(\mathbf{n},z)$ $=$ $27\times 10^{-3}\hskip 2.84544pt\frac{x_{\mathrm{HI}}(\mathbf{n},z)}{T_{S}(\mathbf{n},z)}\left[1+\delta_{\mathrm{b}}(\mathbf{n},z)\right]\left(\frac{\Omega_{\mathrm{b}}h^{2}}{0.023}\right)$ (2)
$\times$ $\left(\frac{0.15}{\Omega_{\mathrm{m}}h^{2}\cdot 10}\right)^{1/2}(1+z)^{3/2}\,$,
where $x_{\rm HI}(\textbf{n},z)$ and $\delta_{\rm b}(\textbf{n},z)$ are neutral hydrogen fraction and baryonic density contrast respectively along the line of sight direction n and redshift $z$. We do not consider the effects of the peculiar velocity in this study.
@@
$\times\left(\frac{0.15}{\Omega_{\rm m}h^{2}}\frac{1+z}{10}\right)^{1/2}\left[1-\frac{T_{\gamma}(z)}{T_{\rm S}(\textbf{n},z)}\right]\,\mathrm{mK}.$ (4)
@@ 2.2 Simulations of Hi  21-cm signal during CD
We use grizzly code (Ghara et al., 2015a, b, 2018) for $21$-cm signal simulations, which is based on a one-dimensional radiative transfer scheme. However, recently Ma et al. (2023); Acharya et al. (2025) developed polar, an algorithm that combines the radiative transfer code grizzly with the semi-analytical galaxy formation code L-Galaxies 2020, providing a self-consistent framework to model galaxy formation, reionization, and the resulting 21-cm signal. grizzly requires a gas density field, a halo catalog and a source model as its inputs. The dark matter distribution in a cubic box of comoving length $714.29$ Mpc is obtained using an N-body code named cubep3m (Harnois-Deraps et al., 2013) which is smoothed into 600 uniform grids resulting in a spatial resolution of $1.19$ Mpc. Dark matter distributions are obtained at multiple redshifts in the range of $z=18$ to $z=10$ during CD. Dark matter halos are identified using an on-the-fly halo finder which uses the spherical overdensity method. The lowest dark matter halos identified in our simulations have masses of approximately $10^{9}$ $\rm M$$_{\odot}$ (the mass of a dark matter particle in our simulations is $\sim 10^{8}$ $\rm M$$_{\odot}$). The stellar mass available within a halo is directly related to the DM halo mass as $M_{*}=f_{*}(\frac{\Omega_{\rm b}}{\Omega_{\rm m}})M_{\rm h}$, where the star formation efficiency $f_{*}$ is assumed to be $0.03$ (Behroozi & Silk, 2015; Sun & Furlanetto, 2016). Our simulation has two different astrophysical parameters i.e., $f_{\rm x}$ and $M_{\rm h,min}$. $M_{\rm h,min}$ is the minimun mass of a dark matter halo capable of hosting sources that emit $\rm{Ly{\alpha}}$, X-ray and UV ionizing photons. We consider two different values of $M_{\rm h,min}$, i.e. $10^{9}$ $\rm M$$_{\odot}$ and $10^{10}$ $\rm M$$_{\odot}$, in this study. $f_{\rm x}$ is the X-ray heating efficiency parameter which determines the X-ray emission rate per unit stellar mass, set as $f_{\rm x}\times 10^{42}\ \mathrm{s}^{-1}\ {\rm M}_{\odot}^{-1}$. We have considered 3 different values of $f_{\rm x}$, i.e. $0.1,46.4$ and $1000$ in our study. The details of these simulations are discussed in Kamran et al. (2022).
We see in eq. 4 that we need to estimate $T_{S}$ in each grid of simulation cube in order to simulate the Hi 21-cm signal i.e., $\delta T_{b}$. The spin temperature $T_{S}$ can be estimated using the equation
$T_{S}^{-1}=\frac{T_{\gamma}^{-1}+x_{\alpha}T_{k}^{-1}}{1+x_{\alpha}},$ (6)
where $T_{k}$ is kinetic temperature of the IGM and $x_{\alpha}$ is the Lyman-$\alpha$ coupling coefficient which can be calculated using the Lyman-$\alpha$ flux $J_{\alpha}$ as,
$x_{\alpha}=\frac{16\pi^{6}T_{*}e^{2}f_{12}J_{0}(|r|)}{27A_{10}T_{k}m_{e}c}.$ (7)
[…]
The 21-cm signal simulations start by solving these one-dimensional coupled differential equations and creating many radial profiles of $x_{\rm HII}$, $T_{k}$ and $x_{\alpha}$ around isolated sources for different combinations of halo masses, UV emissivity, the ratio of X-ray and UV luminosities, redshifts and density contrast. In a post-processing step, these profiles are used to generate the ionization maps and kinetic temperature fields at different redshifts. The related steps for producing H ii maps are estimating and assigning the volume of H ii regions around each source, computing the “unused” ionizing photons in the overlapped regions, and redistributing the unused photons among the overlapping sources by increasing the size of the H ii regions appropriately. The code then computes $x_{\rm HII}$ in the partially ionized regions using a simple overlap prescription. Finally, the kinetic temperature maps are generated using a correlation between $x_{\rm HII}$ and $T_{k}$ (for details, see Ghara et al., 2015a, b, 2018; Islam et al., 2019). The code then generates $T_{s}$ maps using the previously generated $T_{k}$ maps and $x_{\alpha}$ using eq. 6, assuming the collisional coupling is negligible at these redshifts.
@@
Using the DM density distribution obtained from the N-body simulation and spin temperature, neutral hydrogen distribution from the one-dimensional radiative transfer code grizzly, we also generate maps of the Hi 21-cm optical depth $\tau({\bf n},z)$ distribution. We generate $\delta T_{b}$ maps at $23$ different redshifts covering the entire CD using both the approximated (equation (3)) and the exact (equation (1)) $\delta T_{\rm b}$ equations.
@@
We show the evolution of the global signal ($\overline{\delta T_{\rm b}}(z)$) with redshift in the lower panel of Fig. 1. We see, for all choices of $f_{\rm x}$ values and for a fixed $M_{\rm h,min}$, that the initial trend in the evolution of $\overline{\delta T_{\rm b}}(z)$ around redshift $z=18$ is very similar as only $\rm{Ly{\alpha}}$ coupling dictates the 21-cm signal during the initial phase. It helps to couple the spin temperature with the IGM kinetic temperature; therefore, the signal is seen in absorption. Later, X-ray heating becomes important and raises the IGM kinetic and spin temperature. This creates a trough-like feature in the redshift evolution of $\overline{\delta T_{\rm b}}(z)$. Further, we see that the absorption trough becomes deeper for smaller $f_{\rm x}$ values as the IGM heating starts late and the spin temperature, which is coupled to IGM kinetic temperature through $\rm{Ly{\alpha}}$ coupling, continues to decrease due to adiabatic expansion of universe. Moreover, the trough is more flattened for smaller $f_{\rm x}$ values. This is because the X-ray heating rate and cooling rate due to cosmological expansion remain similar for a longer period of time in these scenarios. As expected, higher $f_{\rm x}$ makes $\overline{\delta T_{\rm b}}(z)$ positive at higher redshift. The trough like feature in $\overline{\delta T_{\rm b}}(z)$ shifts towards lower redshifts for higher halo mass cut-off, mainly because massive DM halos are formed later and, therefore, processes such as $\rm{Ly{\alpha}}$ coupling, X-ray heating get delayed.
@@
From the lower panel we also see that if we vary the X-ray heating efficiency $f_{\rm x}$, both the width and depth of the troughs vary. These are maximum for $f_{\rm x}=0.1$ and minimum for $f_{\rm x}=1000$. Further, we see that $\overline{\delta T_{\rm b}}(z)$ calculated using the exact equation is smaller than the approximated method, and the difference is more for smaller values of $f_{\rm x}$. This is expected as $T_{\rm S}$ has lower values for the $f_{\rm x}=0.1$ model, i.e., $\tau$ has higher values, which leads to a greater difference in the troughs between approximated and exact $\overline{\delta T_{\rm b}}(z)$ compared to the $f_{\rm x}=1000$ model. The difference is found to be similar for different $M_{\rm h,min}$ models for a given $f_{\rm x}$ value. The percentage difference between approximated and exact $\overline{\delta T_{\rm b}}(z)$ reaches up to $6.2\%$ for $f_{\rm x}=0.1$, $_{\rm h,min}=10^{9} M{\rm M}_{\odot}$] model and $6.7\%$ for $f_{\rm x}=0.1$, $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$] model. The differences reduce to roughly $5\%$ and $2.7\%$ for $f_{\rm x}=46.4$ and $1000$, respectively, for both the $M_{\rm h,min}$ values. This difference arises because of considerably large values of $\tau$ during the CD. The higher-order terms in $\exp(-\tau)$ become significant when $\tau$ is large. To elaborate on this, we plot the evolution of the mean $\tau$ with redshift in the upper panel of Fig. 1. The redshift at which the mean $\tau$ reaches its maximum corresponds to the largest difference between the exact and approximated $\overline{\delta T_{\rm b}}(z)$. The optical depth becomes maximum when the IGM kinetic and spin temperature are minimum, which occurs just before the onset of X-ray heating.